OpenStudy (anonymous):

A special window has the shape of a rectangle surmounted by an equilateral triangle. If the perimeter of the window is 16 ft, what dimensions will admit the most light?

OpenStudy (anonymous):



OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

i dont know where to start..if you get me started with an equation i can do this : )

hartnn (hartnn):

whats the criteria for admitting most light ?

hartnn (hartnn):

like max. area or something ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what dimensions give the max area is what this is asking

hartnn (hartnn):

so let the length be = x what is width = ? from perimeter formula....

OpenStudy (anonymous):

If x is the length of the rectangle (and side of the triangle), let height of rectangle be h. Then the perimeter of the figure is 3x+2h=16 =>h=8-3x/2 The area of the figure, A, is given by A=xh+(sqrt(3)/4)x^2 =x(8-3x/2)+(sqrt(3)/4)x^2 Differentiating gives dA/dx=8-3x+(sqrt(3)/2)x d^2A/dx^2=-3+(sqrt(3)/2)<0 =>Maximum So maximum area of window is given when 8-x(3-(sqrt(3)/2))=0 =>x=3.75 (2dp) =>h=2.38 (2dp) ie the window is of length 3.75 feet and width 2.38 feet, and the equilateral triangle has sides of 3.75 feet.

OpenStudy (anonymous):

im confused...a lot

OpenStudy (anonymous):

the perimeter formula for the square will be 2x+2y=16

hartnn (hartnn):

but thats the perimeter of entire window, 16 ft

hartnn (hartnn):